OpenStudy (anonymous):

A student has the following grades: 86, 80, 79, 85, 70. Find the standard deviation of the grades.

OpenStudy (amistre64):

sum (x-mean)^2 -------------- n-1 and the sqrt of that

OpenStudy (amistre64):

or just list it in your ti83 and let it do the work

OpenStudy (anonymous):

k

OpenStudy (amistre64):

mean = 80 sd = 6.364 rounded

OpenStudy (anonymous):

dang your a fast worker i struggle in math cause i jst dnt understand and get headaches from it

OpenStudy (amistre64):

86-80 = 6, 36 80-80 = 0, 0 79-80 = 1, 1 85-80 = - 5, 25 70-80 = -10, 100 ------- 162 sqrt(162/(5-1)) sqrt(162/4) sqrt(162)/2 = 6.364 rounded :)

OpenStudy (amistre64):

\[\frac{sum.of(x-mean)^2}{(number.of.data)-1}\]